OpenStudy (anonymous):

Help,fast Please. Thanks for v=<2,-3> ,what is the graph of 2v

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

same vector but double length

OpenStudy (whpalmer4):

yes, to multiply a scalar (such as 2) with a vector (such as v=<2,-3>) you just multiply the scalar by each component of the vector and stuff the result back in the same spot. v=<1,4,7> 3v = <3*1,3*4,3*7> = <3,12,21> To add or subtract, you just add or subtract the corresponding components: v = <1,2,3> u = <2,0,-1> v+u = <1+2,2+0,3-1> = <3,2,2>